# Using Symbolic Regression to Emulate the Radial Fourier Transform of the Sérsic profile for Fast, Accurate and Differentiable Galaxy Profile Fitting

### v1

@@ 4.1 Injection-Recovery Tests
Our final equation is selected based on fitting to the radial Fourier Transform of the Sérsic profile; however, this does not necessarily translate into accuracy in the recovered properties of interest when fitting a galaxy (e.g., $n$, $r_{e}$). To ensure our chosen equation is accurate in real space, we perform a set of injection recovery tests. To begin, we implement this emulator for rendering in pysersic. [^2] This is currently implemented in the branch sersic_fourier_emulator here and will be merged with the main branch and marked with a new version and release upon acceptance of this paper. Given the linearity of the Hankel transform, it is straightforward to scale the emulator $\tilde{\mathcal{F}_{r}}$$\leavevmode\nobreak\ (k,n)$, which assumes a flux and radii of unity, to any arbitrary values during the fitting process. Our “true” profile will be rendered using a heavily over-sampled pixel space renderer. The entire image is over-sampled by a factor of 10 and the inner 8 pixels are over-sampled by an additional factor of 20.
@@
We render a sample of 500 “true” Sérsic profiles onto a 200 pixel by 200 pixel image with $r_{\rm eff}$ randomly sampled between 1-10 pix, and $n$ between 0.56.00.5-6.0. We additionally sample random values of the axis ratio, position angle and central position, the flux of each profile is assumed to be 100. We then find the best fit to each profile using the Fourier Emulator rendering scheme to ensure we achieve consistent results. To help stabilize the fitting process we add a small amount of Gaussian noise with an r.m.s of $10^{-4}$ per pixel.
The results of our injection recovery tests are displayed in Figure 3. We find strong agreement between the input parameters and those recovered using the Sérsic Fourier Emulator to render the profiles. The average difference between the input and recovered parameters is $0.4\%$ and the scatter is $1.7\%$ for both $r_{\rm eff}$ and $n$. We note that there appears to be one outlier where the effective radius is significantly underestimated. This appears to be an issue with the minimization as the recovered ellipticity and position are offset from the true values and the position angle is stuck at 0, a limit of the prior. These results indicate that our chosen emulator is accurate enough to reproduce Sérsic profiles for use in profile fitting.
Figure 5: Comparing the recovered radius, $R_{\rm eff}$ and Sérsic index,$n$, when using the hybrid and Fourier Emulator rendering methods for a sample of 100 galaxies with HSC imaging. The red line shows the one-to-one relation while the black line shows median calculated using the loess algorithm. (Cappellari et al., 2013). An inset is included for each parameter showcasing a histogram of the fractional differences. We find the two rendering methods provide largely consistent results with the median fractional difference of $1\%$ for both parameters with a scatter of $2.6\%$ for $r_{\rm eff.}$ and $4.9\%$ for $n$.
@@ 4.2 Real Galaxies
To test the performance of this method in a realistic scenario we select 100 galaxies from the Galaxy Mass And Assembly (GAMA) survey, Data Release 4, at $0.4<z<0.5$ in the G02 field  (Driver et al., 2022). This field overlaps with the HSC-Subaru Strategic Program imaging program which we will use to perform the morphological fits. Focusing on the $i$ band, science cutouts, weight maps and PSFs are downloaded from the public data release 2 using the unagi python package [^3] https://github.com/dr-guangtou/unagi/tree/master (Aihara et al., 2019). Our aim is to assess the accuracy and performance gain of the rendering method based on the Sérsic Fourier emulator compared to the default hybrid method in pysersic.
@@
We perform inference using two methods. The first is stochastic variational inference (SVI) (Hoffman et al., 2013). This method approximates the posterior distribution using a normalizing flow (De Cao et al., 2020). We also use a minimizer to estimate the maximum-a-posteriori (MAP) for each galaxy. It is generally faster but can be less accurate as it will only ever approximate the posterior. The next is MCMC sampling using the No U-Turn sampler implemented in numpyro (Hoffman et al., 2014; Phan et al., 2019) using 2 chains with $1,000$ warm-up and $1,000$ sampling steps. Before sampling we reparameterize the parameters using a multi-variate normal following  Hoffman et al. (2019). For these tests we use the default settings in pysersic. These tests are all performed on a 2023 Macbook Pro with an M2 Pro processor.
@@
A quantitative comparison of the recovered parameters using the different methods is displayed in Figure 5. Similar to Fig. 3 the recovered half-light radius and Sérsic index from the two different rendering methods are directly compared and the inset histogram shows the distribution of fractional differences, denoted as $\Delta$. We find good agreement between the two methods with the mean fractional difference, calculated using the biweight location (Beers et al., 1990), to be $1\%$ for both parameters. The scatter in $\Delta$, calculated using the biweight scale, is $2.6\%$ for $r_{\rm eff.}$ and $4.9\%$ for $n$. There are minor systematic differences for individual galaxies but the two rendering methods produce very similar results.
@@ 5 Summary and Discussion
Galaxy profile fitting underpins much of the science performed with extragalactic surveys. The most common parameterization is the Sérsic  profile, which can reasonably represent most galaxies but is difficult to work with analytically, including that there is not a known closed form Fourier transform. In this paper we have explored using symbolic regression to emulate the radial Fourier transform, or Hankel transform, of the Sérsic profile. Our goal is to implement an accurate, computationally efficient, and differentiable rendering scheme for galaxy profile fitting.
To achieve this goal we numerically calculated $\mathcal{F}_{r}$$(k,n)$ for a set of $k$, the spatial frequency coordinate and $n$ the Sérsic  index, values using the hankel python package. (Murray & Poulin, 2019). Analyzing the results of our symbolic regression fitting, performed using pysr (Cranmer, 2023), in the context of profile fitting, we selected an equation that is both accurate and computationally efficient. The results of the symbolic regression fitting are shown in Figure 6 and selection equation in Equation 6.
To ensure this emulator is accurate enough to be used in real profile fitting, we performed a set of recovery tests using a heavily over-sampled pixel rendering method as the ‘truth’, and found consistent results, with the average fractional difference in the inferred effective radius and Sérsic index to be $0.4\%$ for both parameters with scatter $1.7\%$.
Finally, to test this renderer in a realistic scenario, we measured the morphological properties of 100 galaxies in the HSC-SSP imaging. Using pysersic, we fit the same set of galaxies using the Sérsic Fourier Emulator rendering method and the default hybrid method based on approximating the Sérsic profile using a mixture of Gaussians. Similar to the injection-recovery tests we found consistent results when comparing the two rendering methods.
@@
Additionally, this approach is not without its limitations. First, it is an empirical approximation, and therefore is sensitive to inaccuracies and only valid for the range of parameters it has been trained on. The range of $n$ values we have used, $0.5<n<6$ encompasses the majority of real galaxies  (e.g. Lange et al., 2015; Mowla et al., 2019) but it is common to see a higher $n$ limit, often up to $n=8$, used in profile fitting. Initially we attempted to find an emulator over the entire range of $0.5<n<8$ however we found the symbolic regression procedure had difficulty emulating the behavior of $\mathcal{F}_{r}$ $(k,n)$ for values of $n>6$. To ensure an accurate equation which covers the vast majority of realistic galaxy profiles we therefore limited our training set, and thus the applicability of our emulator, to profiles with $n<6$.
Another is aliasing, or “wrapping around” of extended sources near the edge of the image. This is a generic issue of rendering in Fourier space due to the periodic nature of the inverse FFT. Lang (2020) introduces a method to overcome this issue using a series of Gaussian to represent the Sérsic profile and render the largest components in real space. This avoids rendering most of the extended emission, which is the most susceptible to aliasing, in Fourier space therefore minimizing the effects of aliasing. Another option is to render on a larger grid and trim the edges back to the original shape. This simplistic approach simply removes the edges of the cutout where the aliasing is most prominent. However, doing so adds significant computational cost as larger grids need to be used.

### v2

@@ 4.1 Injection-Recovery Tests
Our final equation is selected based on fitting to the radial Fourier Transform of the Sérsic profile; however, this does not necessarily translate into accuracy in the recovered properties of interest when fitting a galaxy (e.g., $n$, $r_{e}$). To ensure our chosen equation is accurate in real space, we perform a set of injection recovery tests using a “perfect” Sérsic profile.
To begin, we implement this emulator for rendering in pysersic. [^3] This is currently implemented in the branch sersic_fourier_emulator here and will be included in the next release of v0.2 Given the linearity of the Hankel transform, it is straightforward to scale the emulator $\tilde{\mathcal{F}_{r}}$$\penalty 10000\ (k,n)$, which assumes a flux and radii of unity, to any arbitrary values during the fitting process. Our “true” profile will be rendered using a heavily over-sampled pixel space renderer. The entire image is over-sampled by a factor of 10 and the inner 8 pixels are over-sampled by an additional factor of 20.
Figure 5: Residuals from the injection recovery test for $r_{\rm eff}$ and $n$ as a function of the input value for each parameter. This figure showcases the same data as shown in Fig. 4. The residuals for $r_{\rm eff}$ do not vary systematically however for $n$ we find the biggest disagreement at around $n\lesssim 1.5$ mirroring the results shown in Fig. 3.
@@
We render a sample of 500 “true” Sérsic profiles onto a 200 pixel by 200 pixel image with $r_{\rm eff}$ randomly sampled between 1-10 pix, and $n$ between 0.56.00.5-6.0. We additionally sample random values of the axis ratio uniformly between 0 and 0.9, position angle uniformly between 0 and $\pi$ and central position in each coordinate following a Gaussian centered on the center of the image with a width of 0.5 pixels. The flux of each profile is assumed to be 100. We then find the best fit to each profile using the Fourier Emulator rendering scheme to ensure we achieve consistent results. To help stabilize the fitting process we add a small amount of Gaussian noise with an r.m.s of $10^{-4}$ per pixel.
We note that in our initial optimization there was one source which was an outlier from the distribution, with best fit parameters $20\%$ different from the input values. Upon further investigation we found that the best-fit position angle was exactly 0, the limit of our prior, when the input was $2.66$. This suggested an issue with the minimization. To confirm this we restarted the optimization using a different random seed and found the best-fit parameters were in much better alignment with the input values. We have chosen to only show this second minimization for this source and omit the initial run.
The results of our injection recovery tests are displayed in Figure 4. We find strong agreement between the input parameters and those recovered using the Sérsic Fourier Emulator to render the profiles. For the central position, $x_{c}$ and $y_{c}$, ellipticity and position angle we find agreement to a level better than $10^{-3}$ The average difference between the input and recovered parameters is $0.4\%$ and the scatter is $1.7\%$ for $r_{\rm eff}$ and $n$ and $0.5\%$ for the total flux.
To further investigate the accuracy of the new rendering scheme we show the residuals of the recovery test for $n$ and $r_{\rm eff}$ as a function of the input values. The $r_{\rm eff}$ residuals do not show much structure, with the distribution around the true value remaining relatively consistent as a function of input radius. The residuals of $n$, do show systematic variation as a function of the input index. Most notably the emulator systematically underestimates the index by up to $5\%$ near $n=1$. For values of $n>3$ the index is instead overestimated by $1\%$. This mirrors the results shown in Fig. 3 which indicated the emulator was less accurate for profiles with $n\lesssim 1$.
Figure 7: Comparing the recovered radius, $R_{\rm eff}$ and Sérsic index,$n$, when using the hybrid and Fourier Emulator rendering methods for a sample of 100 galaxies with HSC imaging. The red line shows the one-to-one relation. Two insets are included for each parameter showcasing a histogram of the fractional difference in the upper left, and difference with respect to the uncertainties in the bottom right. For the latter the black dotted line showcases a standard normal distribution, which would be expected if there are no systematic differences between the two. We find the two rendering methods provide largely consistent results with the median fractional difference of $1\%$ for both parameters with a scatter of $2.6\%$ for $r_{\rm eff.}$ and $4.9\%$ for $n$ along with the $\sigma$ differences largely following a standard normal distribution, with some outlier notably in $n$.
@@ 4.2 Real Galaxies
Along with the injection-recovery tests we also wish to test how well this rendering technique performs in a realistic scenario. For this we compare to the hybrid method implemented in pysersic, which is the default. However both of these rendering methods are imperfect. Our goal with this test is not to assess absolute accuracy, as with the injection-recovery tests, but ensure we are able to provide consistent results with an established technique.
We select 100 galaxies from the Galaxy Mass And Assembly (GAMA) survey, Data Release 4, at $0.4<z<0.5$ in the G02 field (Driver et al., 2022). This field overlaps with the HSC-Subaru Strategic Program imaging program which we will use to perform the morphological fits. Focusing on the $i$ band, science cutouts, weight maps and PSFs are downloaded from the public data release 2 using the unagi python package [^4] https://github.com/dr-guangtou/unagi/tree/master (Aihara et al., 2019). Our aim is to assess the accuracy and performance gain of the rendering method based on the Sérsic Fourier emulator compared to the default hybrid method in pysersic.
@@
We perform inference using two methods. The first is stochastic variational inference (SVI) (Hoffman et al., 2013). This method approximates the posterior distribution using a normalizing flow (De Cao et al., 2020). We also use a minimizer to estimate the maximum-a-posteriori (MAP) for each galaxy. It is generally faster but can be less accurate as it will only ever approximate the posterior. The next is MCMC sampling using the No U-Turn sampler implemented in numpyro (Hoffman et al., 2014; Phan et al., 2019) using 2 chains with $1,000$ warm-up and $1,000$ sampling steps. Before sampling we reparameterize the parameters using a multi-variate normal following Hoffman et al. (2019). For these tests we use the default settings in pysersic. We ensure the resulting chains all have effective sample sizes >300>300 for all parameters and $\hat{r}<1.05$ for each renderer (Vehtari et al., 2021). This removes 8 galaxies. Of the remaining galaxies the median effective sample size is 1148 and all values of $\hat{r}$ are $\leq 1.01$. These tests are all performed on a 2023 Macbook Pro with an M2 Pro processor utilizing the CPU.
@@
A quantitative comparison of the recovered parameters using the different methods is displayed in Figure 7. Similar to Fig. 4 the recovered half-light radius and Sérsic index from the two different rendering methods are directly compared. There are two inset histograms: The upper left the distribution of fractional differences, denoted as $\Delta_{f}$ and the bottom right showcasing the difference with respect to the combined uncertainty, denoted $\Delta_{\sigma}$. The combined uncertainty is calculated by adding that measured from each method in quadrature, assuming they are independent.
We find good agreement between the two methods with the mean fractional difference, calculated using the biweight location (Beers et al., 1990), to be $1\%$ for both parameters. The scatter in $\Delta_{F}$, calculated using the biweight scale, is $2.6\%$ for $r_{\rm eff.}$ and $4.9\%$ for $n$. We compare the distribution $\Delta_{\sigma}$ to a standard normal which one would expect to find if there were no systematic differences between the two. For $r_{\rm eff}$ we find that $\Delta_{\sigma}$ largely follows a normal distribution with a slight bias towards negative values. For $n$, the distribution of $\Delta_{\sigma}$ shows outliers at $5\sigma$. These appear to be concentrated at small $n$, near 1, mirroring the results shown in Figure 3. The distribution of $\Delta_{\sigma}$ for $n$ is centered at 0, with a width that appears to be slightly narrower than 1.
@@ 5 Summary and Discussion
Galaxy profile fitting underpins much of the science performed with extragalactic surveys. The most common parameterization is the Sérsic profile, which can reasonably represent most galaxies but is difficult to work with analytically. Crucially, there is not a general known closed form Fourier transform. In this paper we have explored using symbolic regression to emulate the radial Fourier transform, or Hankel transform, of the Sérsic profile. Our goal is to implement an accurate, computationally efficient, and differentiable rendering scheme for galaxy profile fitting.
To achieve this goal we numerically calculated $\mathcal{F}_{r}$$(k,n)$ for a set of $k$, the spatial frequency coordinate and $n$ the Sérsic index, values using the hankel python package. (Murray and Poulin, 2019). Analyzing the results of our symbolic regression fitting, performed using pysr (Cranmer, 2023), in the context of profile fitting, we selected an equation that is both accurate and computationally efficient. The results of the symbolic regression fitting are shown in Figure 2 and the selected equation in Equation 6.
To test this new rendering method we perform two sets of tests. The first is a set of injection-recovery tests using a heavily over-sampled pixel rendering method as a “perfect” profile. We find the emulator provides consistent results with the average fractional difference in the inferred effective radius and Sérsic index to be $0.4\%$ for both parameters with scatter $1.7\%$. However there is systematic variation in the accuracy of the index, with a $5\%$ offset at $n\approx 1$ and $2\%$ offset at $n>3$. The second is comparing this rendering method to the hybrid method implemented in pysersic applied to the same set of 100 galaxies in the HSC-SSP imaging. This is meant to test how it compares to established techniques in a real-world scenario. Using pysersic, we fit the same set of galaxies using the Sérsic Fourier Emulator rendering method and the default hybrid method based on approximating the Sérsic profile using a mixture of Gaussians. Similar to the injection-recovery tests we found consistent results when comparing the two rendering methods.
The systematic uncertainties imparted by the rendering methods proposed here are below the typical uncertainty of morphological catalogs. In order to assess the typical uncertainty we download two publicly available catalogs from van der Wel et al. (2012); George et al. (2024). The van der Wel et al. (2012) catalog is based on HST data and shows a median fractional uncertainty of 8.9% in the radius and 21.5% in the index. George et al. (2024) uses data from the CLAUDS and HSC-SSP surveys and find a median fractional uncertainties of 7.5% for radius and 11.0% for the index. These are all much greater than the systematic uncertainty implied by our injection-recovery tests in Fig. 4. On a catalog level, the level of accuracy the emulator provides appears to be sufficient, however specific science goals may require more precision.
@@
[…]
Another is aliasing, or “wrapping around” of extended sources near the edge of the image. This is a generic issue of rendering in Fourier space due to the periodic nature of the inverse FFT. Lang (2020) introduces a method to overcome this issue using a series of Gaussians to represent the Sérsic profile and render the largest components in real space. This avoids rendering most of the extended emission, which is the most susceptible to aliasing, in Fourier space therefore minimizing the effects of aliasing. Another option is to render on a larger grid and trim the edges back to the original shape. This simplistic approach simply removes the edges of the cutout where the aliasing is most prominent. However, doing so adds significant computational cost as larger grids need to be used.